OpenStudy (tylerascough):

what is 1/8 of 48 will fan and medal!!!!!

OpenStudy (tylerascough):

Anyone

Directrix (directrix):

1/8 of 48 means 1/8 * 48 1/8 * 48 = 1/8 * 48/1 = 48/8 = 24/4 = 6 -------- 1/8 of 48 = 6 @TylerAscough